please answer all parts and show work.

20) The equilibrium constant for e reaction Ni2+ (
Answer #1

Part A)

Using the following equation we get the dG

\Delta G^{0}=-RTln(K)

\Delta G^{0}=-(8.314\frac{Pam3}{molK})(298K)ln(5.6x10^{8})

\Delta G^{0}= -49,906.8 J/mol

Part B

As dG<<0, the process is spontaneous, therefore the reaction at standart conditions of P and T is product favored.

Part C

We need to get K

K=\frac{[Ni(NH3)6]}{[Ni][NH3]^{6}}

K=\frac{(0.010)}{(0.0010)(0.0050)^{6}}=6.4X10^{14}

Now using the equation [1] we find dG

\Delta G^{0}=-(8.314)(298K)ln(6.4x10^{14})

\Delta G^{0}= -84,466.6 J/mol

Part D

I cant' see the sketch. Please upload a photo with that part of the problem.